# In deep reinforcement learning, a pruned network is a good network

### v1

@@ 5 Why is pruning so effective?
We focus our analyses on four games: BeamRider, Breakout, Enduro, and VideoPinball. For each, we measure the variance of the $Q$ estimates (QVariance); the average norm of the network parameters (ParametersNorm); the average norm of the $Q$-values (QNorm); the effective rank of the matrix (Srank) as suggested by Kumar et al. (2021a), and the fraction of dormant neurons as defined by Sokar et al. (2023). We present our results in Figure 11. What becomes evident from these figures is that pruning (i) reduces variance, (ii) reduces the norms of the parameters, (iii) decreases the number of dormant neurons, and (iv) increases the effective rank of the parameters. Some of these observations can be attributed to a form of normalization, whereas others may arise due to increased network plasticity.
@@ 5.1 Comparison to other methods
We compare against two recent works that proposed methods for directly dealing with loss of plasticity. Nikishin et al. (2022) observed a decline in performance with an increased replay ratio, attributing it to overfitting on early samples, an effect they termed the “primacy bias”. The authors suggested periodically resetting the network and demonstrated that it proved very effective at mitigating the primacy bias, and overfitting in general (this is labeled as Reset in our results). Sokar et al. (2023) demonstrated that most deep RL agents suffer from the dormant neuron phenomenon, whereby neurons increasingly “turn off” during training of deep RL agents, thus reducing network expressivity. To mitigate this, they proposed a simple and effective method that Recycles Dormant neurons (ReDo) throughout training.
As Figure 12 illustrates, gradual magnitude pruning surpasses all the other regularization methods at all levels of scale, and throughout the entirety of training. Interestingly, most of the regularization methods suffer a degradation when increasing network width. This suggests that the effect of pruning cannot be solely attributed to either a form of normalization or plasticity injection. However, as we will see below, increased plasticity does seem to arise out of its use.
@@ 6 Discussion and Conclusion
As RL continues to grow outside of academic benchmarks and into more complex tasks, it is almost surely going to necessitate larger, and more expressive, networks. In this case parameter efficiency becomes crucial to avoid the performance collapse prior works have shown, as well as for reducing computational costs (Ceron and Castro, 2021).
Our work provides convincing evidence that sparse training techniques such as gradual magnitude pruning can be effective at maximizing network utilization. Figures 1 and 4 suggest a type of “scaling law”, the likes of which has mostly eluded deep RL networks. The results in Figures 6, 8, and 10 all demonstrate that the sparse networks produced by pruning are better at maintaining stable performance when trained for longer.
@@
It would be natural, then to explore incorporating gradual magnitude pruning into recent agents that were designed for multi-task generalization (Taiga et al., 2023; Kumar et al., 2022), sample efficiency (Schwarzer et al., 2023; D’Oro et al., 2023), and generalizability (Hafner et al., 2023). Further, the observed stability of the pruned networks may have implications for methods which rely on fine-tuning or reincarnation (Agarwal et al., 2022).
@@
At a minimum, we hope this work serves as an invitation to explore non-standard network architectures and topologies as an effective mechanism for maximizing the performance of reinforcement learning agents.
@@ Acknowledgements
The authors would like to thank Laura Graesser, Gopeshh Subbaraj and the rest of the Google DeepMind Montreal team for valuable discussions during the preparation of this work. We would also like to thank the Python community Van Rossum and Drake Jr (1995); Oliphant (2007) for developing tools that enabled this work, including NumPy Harris et al. (2020), Matplotlib Hunter (2007) and JAX Bradbury et al. (2018).
@@ Appendix B Atari Game Selection
15 game subset: MsPacman, Pong, Qbert, (Assault, Asterix, BeamRider, Boxing, Breakout, CrazyClimber, DemonAttack, Enduro, FishingDerby, SpaceInvaders, Tutankham, VideoPinball. According to Graesser et al. (2022), these games were selected to be roughly evenly distributed amongst the games ranked by DQN’s human normalized score in (Mnih et al., 2015) with a lower cut off of approximately 100% of human performance.
@@ Appendix C Sparsity Levels
height 90pt depth 0 pt width 1.2 pt Refer to captionRefer to caption
@@
Figure 15: Evaluating how varying the sparsity parameter affects performance for a given architecture, resnet with width=3, on Rainbow agent. We report results aggregated IQM of human-normalized scores over 15 games.
@@ Appendix D Scaling Replay Ratios
Figure 16: Scaling replay ratio for resnet architecture (default is 0.25), for width=$1$ (left) and width=$3$ (right) using DQN agent. We report interquantile mean performance with error bars indicating 95% confidence intervals. On the x-axis we report the replay ratio value.
@@ Appendix E MuJoCo environments
Figure 17: Evaluating how varying the sparsity parameter affects performance of SAC on three MuJoCo environments when increasing width x5. We report returns over 10 runs for each experiment.

### v3

@@ 5 Why is pruning so effective?
We focus our analyses on four games: BeamRider, Breakout, Enduro, and VideoPinball. For each, we measure the variance of the $Q$ estimates (QVariance); the average norm of the network parameters (ParametersNorm); the average norm of the $Q$-values (QNorm); the effective rank of the matrix (Srank) as suggested by Kumar et al. (2021a), and the fraction of dormant neurons as defined by Sokar et al. (2023).
We present our results in Figure 11. What becomes evident from these figures is that pruning (i) reduces variance, (ii) reduces the norms of the parameters, (iii) decreases the number of dormant neurons, and (iv) increases the effective rank of the parameters. Some of these observations can be attributed to a form of normalization, whereas others may arise due to increased network plasticity.
Lyle et al. (2024) show that increased parameter norm accompanies plasticity loss in different neural architectures. In Figure 11, we observe a low parameter norm value when applying gradual pruning, which represent a high final performance return.
@@ 5.1 Comparison to other methods
Lottery ticket baseline
Frankle & Carbin (2018) argued that neural networks contain sparse sub-networks that can be trained at high levels of sparsity without gradual pruning; the authors provide an algorithm for finding these winning tickets. After training a network with pruning, we train a new network with the final mask fixed (i.e. not adjusted during training) and with the parameters initialized as in the original dense network. We found that the proposal under-performs both the pruning approach and the unpruned baseline. It is interesting to observe that both the pruning approach and the lottery ticket experiment seem to still be progressing at 40M, whereas the baseline seems to start deteriorating (Figure 13).
Figure 13: Lottery ticket hypothesis experiment. Taking the final pruned network (with a width factor of 3) and retraining with the original initialization results in worse performance.
Dynamic sparse training baselines
Evci et al. (2020) proposed RigL as a dynamic sparse training mechanism that maintains a sparse network throughout the entirety of training. In Section G.3 we evaluated RigL at various sparsity levels and found that, while effective, RigL is unable to match the performance of gradual magnitude pruning; these results are consistent with those of Graesser et al. (2022).
@@
We compare against two recent works that proposed methods for directly dealing with loss of plasticity. Nikishin et al. (2022) observed a decline in performance with an increased replay ratio, attributing it to overfitting on early samples, an effect they termed the “primacy bias”. The authors suggested periodically resetting the network and demonstrated that it proved very effective at mitigating the primacy bias, and overfitting in general (this is labeled as Reset in our results).
Sokar et al. (2023) demonstrated that most deep RL agents suffer from the dormant neuron phenomenon, whereby neurons increasingly “turn off” during training of deep RL agents, thus reducing network expressivity. To mitigate this, they proposed a simple and effective method that Recycles Dormant neurons (ReDo) throughout training.
As Figure 13 illustrates, gradual magnitude pruning surpasses all the other regularization methods at all levels of scale, and throughout the entirety of training. Interestingly, most of the regularization methods suffer a degradation when increasing network width. This suggests that the effect of pruning cannot be solely attributed to either a form of normalization or plasticity injection. However, as we will see below, increased plasticity does seem to arise out of its use. We provide sweeps over various baseline hyperparameters in Sections G.5, G.6, G.7 and G.8.
@@ 6 Discussion and Conclusion
As RL continues to grow outside of academic benchmarks and into more complex tasks, it is almost surely going to necessitate larger, and more expressive, networks. In this case parameter efficiency becomes crucial to avoid the performance collapse prior works have shown, as well as for reducing computational costs (Ceron & Castro, 2021).
Our work provides convincing evidence that sparse training techniques such as gradual magnitude pruning can be effective at maximizing network utilization, especially as the initial networks are scaled (see Figures 1 and 4). The results in Figures 8, 7, and 10 all demonstrate that the sparse networks produced by pruning are better at maintaining stable performance when trained for longer. The advantages of pruning remain even when sweeping over various baseline hyper-parameters (see Sections G.9, G.4, G.11 and G.10). It is worth noting that the performance of the dense baselines does improve when adjusting the learning rate based on the width multiplier (Section G.9); however, pruning is still the most performant in these settings.
@@
Future work
It would be natural to explore incorporating gradual magnitude pruning into recent agents that were designed for multi-task generalization (Taiga et al., 2023; Kumar et al., 2022), sample efficiency (Schwarzer et al., 2023; D’Oro et al., 2023), and generalizability (Hafner et al., 2023). Further, the observed stability of the pruned networks may have implications for methods which rely on fine-tuning or reincarnation (Agarwal et al., 2022).
@@
At a minimum, we hope this work serves as an invitation to explore non-standard network architectures and topologies as an effective mechanism for maximizing the performance of reinforcement learning agents. Reinforcement learning agents typically use networks originally designed for stationary problems; therefore, other topologies might be better suited to the non-stationary nature of RL.
@@ Acknowledgements
The authors would like to thank Laura Graesser, Utku Evci, Gopeshh Subbaraj, Evgenii Nikishin, Hugo Larochelle, Ayoub Echchahed, Zhixuan Lin and the rest of the Google DeepMind Montreal team for valuable discussions during the preparation of this work.
Laura Graesser deserves a special mention for providing us valuable feed-back on an early draft of the paper. We thank the anonymous reviewers for their valuable help in improving our manuscript. We would also like to thank the Python community (Van Rossum & Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), Jupyter (Kluyver et al., 2016), Pandas (McKinney, 2013) and JAX (Bradbury et al., 2018).
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning, and reinforcement learning in particular. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Atari Game Selection
15 game subset: MsPacman, Pong, Qbert, (Assault, Asterix, BeamRider, Boxing, Breakout, CrazyClimber, DemonAttack, Enduro, FishingDerby, SpaceInvaders, Tutankham, VideoPinball. According to (Graesser et al., 2022), these games were selected to be roughly evenly distributed amongst the games ranked by DQN’s human normalized score in (Mnih et al., 2015) with a lower cut off of approximately $100\%$ of human performance.
@@ Appendix C Sparsity Levels
height 100pt depth 0 pt width 1.2 pt Refer to captionRefer to caption
@@
Figure 16: Evaluating how varying the sparsity parameter affects performance for a given architecture, resnet with a width multiplier of $3$, on Rainbow agent. We report results aggregated IQM of human-normalized scores over $15$ games.
@@ Appendix D Scaling Replay Ratios
Figure 17: Scaling replay ratio for resnet architecture (default is $0.25$), for a width factor of $1$ (left) and a width factor of $3$ (right) using DQN agent. We report interquantile mean performance with error bars indicating 95% confidence intervals. On the x-axis we report the replay ratio value.
@@ Appendix E MuJoCo environments
Figure 18: Evaluating how varying the sparsity parameter affects performance of SAC on three MuJoCo environments when increasing width x5555. We report returns over 10 runs for each experiment.
Appendix F Hyper-parameters list
Default hyper-parameter settings for DER (Van Hasselt et al., 2019) and DrQ($\epsilon$) (Kaiser et al., 2020; Agarwal et al., 2021). Table 1 shows the default values for each hyper-parameter across all the Atari games.
Table 1: Default hyper-parameters setting for DER and DrQ($\epsilon$) agents.
Atari
Hyper-parameter DER DrQ($\epsilon$)
Adam’s($\epsilon$) 0.00015 0.00015
Adam’s learning rate 0.0001 0.0001
Batch Size 32 32
Conv. Activation Function ReLU ReLU
Convolutional Width 1 1
Dense Activation Function ReLU ReLU
Dense Width 512 512
Normalization None None
Discount Factor 0.99 0.99
Exploration $\epsilon$ 0.01 0.01
Exploration $\epsilon$ decay 2000 5000
Minimum Replay History 1600 1600
Number of Atoms 51 0
Number of Convolutional Layers 3 3
Number of Dense Layers 2 2
Replay Capacity 1000000 1000000
Reward Clipping True True
Update Horizon 10 10
Update Period 1 1
Weight Decay 0 0
Sticky Actions False False
Default hyper-parameter settings for DQN (Mnih et al., 2015), Rainbow (Hessel et al., 2018), IQN (Dabney et al., 2018), Munchausen-IQN (Vieillard et al., 2020). Table 2 shows the default values for each hyper-parameter across all the Atari games.
Table 2: Default hyper-parameters setting for DQN, Rainbow, IQN, Munchausen-IQN agents.
Atari
Hyper-parameter DQN Rainbow IQN M-IQN
Adam’s ($\epsilon$) 1.5e-4 1.5e-4 3.125e-4 3.125e-4
Adam’s learning rate 6.25e-5 6.25e-5 5e-5 5e-5
Batch Size 32 32 32 32
Conv. Activation Function ReLU ReLU ReLU ReLU
Convolutional Width 1 1 1 1
Dense Activation Function ReLU ReLU ReLU ReLU
Dense Width 512 512 512 512
Normalization None None None None
Discount Factor 0.99 0.99 0.99 0.99
Exploration $\epsilon$ 0.01 0.01 0.01 0.01
Exploration $\epsilon$ decay 250000 250000 250000 250000
Minimum Replay History 20000 20000 20000 20000
Number of Atoms 0 51 - -
Kappa - - 1.0 1.0
Num tau samples - - 64 64
Num tau prime samples - - 64 64
Num quantile samples - - 32 32
Number of Convolutional Layers 3 3 3 3
Number of Dense Layers 2 2 2 2
Replay Capacity 1000000 1000000 1000000 1000000
Reward Clipping True True True True
Update Horizon 1 3 3 3
Update Period 4 4 4 4
Weight Decay 0 0 0 0
Sticky Actions True True True True
Tau - - 0 0.03
Default hyper-parameter settings for CQL (Kumar et al., 2020) and CQL+C51 (Kumar et al., 2022) offline agents. Table 3 shows the default values for each hyper-parameter across all the Atari games.
Table 3: Default hyper-parameters setting for CQL and CQL+C51 agents.
Atari
Hyper-parameter CQL CQL+C51
Adam’s($\epsilon$) 0.0003125 0.00015
Batch Size 32 32
Conv. Activation Function ReLU ReLU
Convolutional Width 1 1
Dense Activation Function ReLU ReLU
Normalization None None
Dense Width 512 512
Discount Factor 0.99 0.99
Learning Rate 0.00005 0.0000625
Number of Atoms 0 51
Number of Convolutional Layers 3 3
Number of Dense Layers 2 2
Fixed Replay Capacity 2,500,000 steps 2,500,000 steps
Reward Clipping True True
Update Horizon 1 3
Update Period 1 1
Weight Decay 0 0
Replay Scheme Uniform Uniform
Dueling False True
Double DQN False True
CQL coef 0.1 0.1
Default hyper-parameter settings for CNN architecture (Mnih et al., 2015) and Impala-based ResNet (Espeholt et al., 2018) Table 4 shows the default values for each hyper-parameter across all the Atari games.
Table 4: Default hyper-parameters for neural networks.
Atari
Hyper-parameter CNN architecture (Mnih et al., 2015) Impala-based ResNet (Espeholt et al., 2018)
Observation down-sampling (84, 84) (84, 84)
Frames stacked 4 4
Q-network (channels) 32, 64, 64 32, 64, 64
Q-network (filter size) 8 x 8, 4 x 4, 3 x 3 8 x 8, 4 x 4, 3 x 3
Q-network (stride) 4, 2, 1 4, 2, 1
Num blocks - 2
Use max pooling False True
Skip connections False True
Hardware Tesla P100 GPU Tesla P100 GPU
Appendix G Additional experiments
Unless otherwise specified, in all experiments below we report the interquantile mean after 40 million environment steps, aggregated over 15 games with 5 seeds each; error bars indicate 95% stratified bootstrap confidence intervals (Agarwal et al., 2021).
G.1 Experiments with PPO on MuJoCo
We used the PPO implementation from (Graesser et al., 2022) and ran some initial experiments with MuJoCo, increasing the width by 5x. As with our SAC experiments, we see no real change in performance, with perhaps some mild gains in Humanoid-v2. One reason why we may not see performance improvements in neither SAC nor PPO is that in ALE experiments, all agents use Convolutional layers, whereas for the MuJoCo experiments (where we ran SAC and PPO) the networks only use dense layers.
Nonetheless, it is worth noting that Graesser et al. (2022) saw degradation with pruning at width=1 (Figure 16 in their paper), with almost a total collapse at 99% sparsity. In contrast, our results with 5x width shows strong performance even at 99% sparsity.
Figure 19: Proximal Policy Optimization (PPO) (Schulman et al., 2017) on MuJoCo environments when increasing width x5. We report returns over 10 runs for each experiment.
G.2 Experiments with IQN and M-IQN
While Rainbow is still a competitive agent in the ALE and both DQN and Rainbow are still regularly used as baselines in recent works, exploring newer agents is a reasonable request. To address this, we ran experiments with Implicit Quantile Networks (IQN) (Dabney et al., 2018) and Munchausen-IQN (Vieillard et al., 2020) with widths of 1 and 3; consistent with our submission’s findings, we observe significant gains when using pruning.
Figure 20: IQN (Implicit Quantile Networks) (Dabney et al., 2018) with ResNet architecture (with a width factor of $1$ and $3$).
Figure 21: M-IQN (Munchausen-Implicit Quantile Networks) (Vieillard et al., 2020) with ResNet architecture (with a width factor of $1$ and $3$).
G.3 Comparison with RigL
We have run a comparison with RigL (Evci et al., 2020). While RigL can be somewhat effective, it is unable to match the performance of pruning.
Figure 22: The Rigged Lottery (RigL) (Evci et al., 2020) for DQN with ResNet architecture (with a width factor of $1$ and $3$).
G.4 Varying Adam’s $\epsilon$
The default value for Adam’s $\epsilon$ is $51.5e-$ 5; we ran experiments by dividing/multiplying this value by $3$ ( $55e-$ 5 and $44.5e-$ 4, respectively). In all these cases, pruning maintains a significant advantage over the dense baseline.
Figure 23: Adam’s epsilon ($\epsilon$) (Kingma & Ba, 2014) for DQN with ResNet architecture and a width multiplier of $3$.
G.5 Sweeping over ReDo threshold $\tau$
This parameter (introduced in Definition $3.1$ of (Sokar et al., 2023)) defines the threshold for determining neuron dormancy. Sokar et al. (2023) suggested using 0.1 with the CNN network. Since we are using the Impala network architecture, we tested three additional values: (00, $0.025$, $0.3$). We found that $0.1$, as used in Figure 13 of our submission, yields the best performance.
Figure 24: Varying ReDo’s $\tau$ threshold (Sokar et al., 2023) for DQN with ResNet architecture and a width multiplier of $3$.
G.6 Frequency of network resets
We varied the frequency of network resets (Nikishin et al., 2022) to evaluate whether this could help mitigate the performance loss when increasing the network width. While more infrequent resets (every $250000$ steps compared to the default value of $100000$) improves performance slightly, it still drastically under-performs with respect to the baseline and the pruning approach.
Figure 25: Varying the reset period (Nikishin et al., 2022) for DQN with ResNet architecture and a width multiplier of $3$.
G.7 Layer to reset
In our paper we followed the approach of Nikishin et al. (2022) and Sokar et al. (2023) of resetting only the last layer. We explored resetting different layers, but found it resulted in no significant performance difference.
Figure 26: Resetting different layers (Nikishin et al., 2022), DQN with ResNet architecture and a width multiplier of $3$.
G.8 Varying weight decay
We ran a sweep over the following values $10e^{-6}$ , $10e^{-5}$ , $10e^{-4}$ , $10e^{-3}$, and $10e^{-2}$, using the Impala architecture with a width factor of $3$. The best performance is obtained with $10e^{-5}$, which is the value suggested by Sokar et al. (2023), and the value used in Figure 13 of our submission.
Figure 27: Weight Decay (WD) for DQN with ResNet architecture with a width factor of $3$.
G.9 Varying learning rates
The default learning for DQN is $56.25e-$ 5. As suggested by the reviewer, we have run experiments with a learning rate divided by the width scale factor (so $52.08e-$ 5 for a width factor of $3$, and $51.25e-$ 5 for a width factor of $5$). These learning rates do improve the performance of the baseline, but it is still surpassed by pruning. These results are consistent with the thesis of the paper: pruning can serve as a drop-in mechanism for increasing agent performance.
Figure 28: Learning rate evaluation for DQN agent with ResNet architecture (with a width factor of $3$ left and $5$ right). These learning rate values correspond to dividing the default learning rate by the factor we used to amplify the size of the neural network. The dashed lines indicate the final performance for dense (blue) and $0.95\%$ sparse (orange) nets when using the default learning rate ( lr: $56.25e-$ 5).
G.10 Varying batch size
The default batch size is $32$, and ran experiments with batch sizes of $16$ and $64$. In all cases, pruning maintains its strong advantage.
Figure 29: Batch size (Ceron et al., 2023) for DQN with ResNet architecture and a width multiplier of $3$.
G.11 Varying update horizon
We explored using an update horizon of $3$ for DQN (the default is $1$) and found that pruning still maintains its advantage.
Figure 30: Multi-step return (Sutton, 1988) for DQN with ResNet architecture and a width multiplier of $3$.